Admitting New Partners Who Buy an Interest and Contribute Assets

The capital accounts of Trent Henry and Tim Chou have balances of $163,000 and $117,200, respectively. LeAnne Gilbert and Becky Clarke are to be admitted to the partnership. Gilbert buys one-fifth of Henry’s interest for $37,500 and one-fourth of Chou’s interest for $25,800. Clarke contributes $39,800 cash to the partnership, for which she is to receive an ownership equity of $39,800.

a1. Journalize the entry to record the admission of Gilbert. For a compound transaction, if an amount box does not require an entry, leave it blank.

a2. Journalize the entry to record the admission of Clarke.

b. What are the capital balances of each partner after the admission of the new partners?

Partner Capital Balance
Trent Henry $
Tim Chou $
LeAnne Gilbert $
Becky Clarke $
Answer #1
Particulars Trent Henry Tim chou LeAnne gilbert Becky clarke Total share Capital
Opening capital a/c 163,000 117,200 280,200
Gilbert adj (32,600) (23,440) 56,040
Clarke adj (19,900) (19,900) 39,800
Closing capital a/c 110,500 73,860 56,040 39,800 280200

Journal enteries:-

a1

Bank Alc Dr 57,440

To Gilbert capital A/c 56,040

To Goodwill A/c 1400

(Being capital introduced by the Gilbert)

a2

Bank A/c Dr 39,800

To Clarke capital A/c 39,800

(Being capital introduced by the clarke)

Note: 1 Gilbert Share

1/5 of henry share =163,000/5=32,600,actualy paid 37,500,difference will be the goodwill=37,500-32,600=4900

1/4 of chou share=117,200/4=29,300,actualy paid 25,800,difference will be the goodwill=25,800-29,300=3500

So, overall Goodwill amount= 4900 - 3500 = 1400

2. Clarke Share

Ownership Equity = 39,800

  

b)

Partner Capital Balance
Trent Henry 110,500
Tim Chou 73,860
LeAnne Gilbert 56,040
Becky Clarke 39,800